OpenStudy (anonymous):

please help Asap Determine the best method to solve the system of equations. Then solve the system. 3x-2y-4 2x-2y=-8 A. Elimination using subtraction;(12,16) B. Elimination using addition;(2,6) C. Elimination using subtraction;(2,1) D. Elimination using addition;(-2,-6)

OpenStudy (anonymous):

you unfortunately are missing the equations, maybe a copy and paste fail

OpenStudy (anonymous):

ah there they are

OpenStudy (anonymous):

ya I notice that and I was fixing it

OpenStudy (anonymous):

\[3x-2y=4 \\2x-2y=-8\] still some typos in your equations did i wrote it correctly?

OpenStudy (anonymous):

lol did i WRITE it correctly?

OpenStudy (anonymous):

i guess so i just copy and paste it

OpenStudy (anonymous):

yours was missing a y and an equal sign

OpenStudy (anonymous):

you wrote this :\[3x-2-4\\ 2x-2y=-8\]

OpenStudy (anonymous):

there is a y next to the two -> 3x−2y−4

OpenStudy (anonymous):

ok i am going to imagine it is \[3x-2y=4 \\2x-2y=-8\] then if you subtract the second equation from the first you get \[x=12\] right away

OpenStudy (anonymous):

go with A

OpenStudy (anonymous):

okay and that's what i was thinking at 1st but didn't know for sure